Hi! I really need some help with this problem I've been stuck on.
The length of a rectangle is 8 in. more than its width. The perimeter is 52 in.

What is the length of the rectangle?

Found 2 solutions by solver91311, Cromlix:
Answer by solver91311(24713)



The length is 8 more than the width, so , then since the perimeter is ,



Solve for

John

My calculator said it, I believe it, that settles it

Answer by Cromlix(4381)

Hi there,
Make the width = 'x'
Length = 8 + x
Perimeter = 2 x Width + 2 x Length
52 = 2(x) + 2(8 + x)
52 = 2x + 16 + 2x
Collect like terms.
2x + 2x = 52 - 16
4x = 36
x = 9
Width = 9 ins
Length = 17 ins.
Hope this helps :-)